OpenStudy (anonymous):

What is the value of y in the equation 5x + 2y = 20, when x = 0.3?

OpenStudy (anonymous):

What is the value of y in the equation 5x + 2y = 20, when x = 0.3? 2.5 2.8 9.25 10.75

OpenStudy (anonymous):

9.25

OpenStudy (anonymous):

You just plug in 0.3 as x and solve for y. 5(0.3)+2y=20 1.5+2y=20 2y=18.5 y=9.25